• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name O. Hansen

    Log entry time 16:00:58 on July25,2001

    Entry number 67561

    keyword=Shift Summary

    8:00-10:05 Runs 2249-2250. Finishing data taking at x=0.48, transverse pol. (270 deg), beam lambda/2 OUT.
    10:05 Changing to BeO target. Spectrometers to 1.32 GeV/c.
    10:10 Left arm Q2 quench.
    10:50 Quench cleared. Right arm dipole still settling, so no time lost.
    11:13 Start run 2251/21251 on BeO/carbon, no raster, 5uA, for pointing
    11:50 Start data taking with BeO/12C, raster on, 25uA, false asymmetry checks
    ca. 15:00 Notice much increased value of charge asymmetry (-1789ppm). Apparently the value started rising around 14:00. Bob Michaels investigating.